OpenStudy (anonymous):

Prove sin(8x)sin(2x)=(sin(5x))^2-(sin(3x))^2

OpenStudy (badhi):

$$\begin{align*} \sin(8x)\sin(2x) &= \frac 12 [\cos(6x)-\cos(10x)]\\ &=\frac 12 [1-2\sin^2(3x)-1+2\sin^2(5x)]\\ &=\frac 12[2\sin^2(5x)-2\sin^2(3x)]\\ &=\sin^2(5x)-\sin^2(3x) \end{align*}$$

OpenStudy (anonymous):

Yep thanks

OpenStudy (anonymous):

\[\sin ^{2}A-\sin ^{2}B=\sin \left( A+B \right)\sin \left( A-B \right)\]